At 10am Brad headed west at 8 mph. At noon Danny headed east from the same place at 12mph. What time was it when they were 116 miles apart?
Answer by checkley71(8403)

8x+12(x-2)=116
8x+12x-24=116
20x=116+24
20x=140
x=140/20
x=7 hours after Brad left they will be 116 miles apart.
proof:
8*7+12(7-2)=116
56+12*5=116
56+60=116
116=116